# Towards Generating Executable Metamorphic Relations Using Large Language Models

### v1

@@ IV Experiment results
EXP1 Results. Table I shows the results of EXP1, focusing on the combined number of responses (i.e., Likert ratings) for each statement across the 64 MRs reviewed by the three practitioners at SISW. From the results of statement 1, we found that in 49 out of the 64 MRs (representing 77%), the practitioners were able to understand the presented MRs and expressed positive feedback, i.e., “strongly agree” or “agree”. In 12 MRs, the practitioners were neutral, and in the remaining 3 MRs, they disagreed with statement 1. Regarding statement 2, the practitioners agreed in 41 out of the 64 MRs (64%), that the MRs capture properties that need to be considered when testing the SUT. In 10 MRs, the practitioners were neutral; and in the remaining 13 MRs, the practitioners disagreed or strongly disagreed with statement 2. In contrast to these results, which are more positive, the responses to statement 3 showed mixed ratings. This statement indicates whether the presented MR helps identify the expected outputs for given inputs. In 18 MRs, the practitioners agreed; in 19 MRs, they were neutral; and in the remaining 27, they disagreed or strongly disagreed.
TABLE I: Responses to the MRs survey, for each statement (S1, S2, S3).
strongly agree
agree
neutral
disagree
strongly disagree
S1
3
46
12
3
0
S2
0
41
10
12
1
S3
0
18
19
24
3
@@
Regarding the non-positive responses to statement 3, we identified possible reasons from the practitioners’ qualitative feedback. In particular, one practitioner, who provided 11 ratings as “disagree” and one as “strongly disagree” to the statement, stated that while most of the MRs are correct and useful for reminding someone what to test, they are too generic to aid in deriving appropriate testing. This aligns with feedback from another practitioner at SISW who, despite not assessing the MRs, collected and reviewed the responses to the MRs questionnaire. The practitioner provided a possible reason for some of the MRs being very generic, stating that it could be attributed to the nature of the documents provided by SISW.
@@
Regarding the survey results for the 50 EMRs in EXP1, one practitioner provided consistent responses across all EMRs, as follows: First, the practitioner perceived that the EMRs are clean and understandable. Second, the practitioner found that the EMRs are consistent with their corresponding MRs. However, the practitioner deemed that not all functions used in the EMRs are feasible to be implemented. Regarding the last opinion, the practitioner noted that the functions in the EMRs, generated by ChatGPT, are specific to each EMR. Therefore, there is little opportunity for reusing these functions across different EMRs, meaning the effort required for implementation is significant. In addition, the specific constraints of the testing environment for their applications may pose challenges in using the exposed APIs to implement functions in the EMRs. Even though EXP1 generated incomplete EMRs with regard to using the SUT’s APIs, as discussed above, this finding suggests that the characteristics of the SUT, such as its APIs, should be considered when further elaborating our approach to deriving EMRs.
@@
EXP2 Results. In EXP2, we annotated each line of the ten EMRs, which were converted from the corresponding ten MRs, for the web application studied in the prior work of MST-wi. These EMRs have a minimum of 10 statements, a mean of 13.8 statements, and a maximum of 20 statements, totaling 138 statements.
@@
Out of the 138 statements in the EMR code, the majority (107 statements) were classified as correctly converted. Specifically, 52 of the 107 statements correctly and exactly implement part of the corresponding MRs. 54 of the 107 statements correctly use SMRL constructs. The remaining 1 statement is valid and correct with respect to the corresponding MR. As for the 31 statements that were deemed incorrect, they can be further categorized as follows: (1) 13 statements incorrectly use APIs; (2) 10 statements incorrectly use SMRL constructs;(3) 6 statements incorrectly implement what is suggested in the corresponding explanations; (4) 2 statements miss implementing what described in the corresponding explanations. Note that each EMR statement is associated with an explanation in natural language produced by ChatGPT.
V Conclusion and future work
@@
In this paper, we have introduced our approach for automatically deriving EMRs from requirements using LLMs. In our preliminary experiments, we applied our approach to four SISW applications and conducted a questionnaire-based survey study to collect opinions from SISW practitioners. In addition, we evaluated the correctness of the EMRs generated by our approach, by applying it to a web application.
@@
In our future work, we plan to explore several key areas related to our current research. A primary focus will be on automating prompt engineering to facilitate the generation of prompts for deriving MRs from requirements and their subsequent conversion into EMRs. Recognizing the significance of input quality, particularly in requirements, another important area will involve developing methods to assist engineers in providing high-quality inputs.

### v2

@@ 4 Experiment results
EXP1. Table 6 shows the results of EXP1, focusing on the combined number of responses (i.e., Likert ratings) for each statement across the 64 MRs reviewed by the three practitioners at SISW. From the results of statement S1, we found that in 49 out of the 64 MRs (77%), the practitioners were able to understand the proposed MRs and expressed positive feedback, i.e., “strongly agree” or “agree”. In 12 MRs, the practitioners were neutral, and in the remaining 3 MRs, they disagreed with statement S1. Regarding statement S2, the practitioners agreed that the MRs capture properties that need to be considered when testing the SUT for 41 out of the 64 MRs (64%). In 10 MRs, the practitioners were neutral; in the remaining 13 MRs, the practitioners disagreed or strongly disagreed with statement S2. In contrast to these results, which are more positive, the responses to statement S3 showed mixed ratings. This statement indicates whether the presented MR helps identify the expected outputs for given inputs. In 18 MRs, the practitioners agreed; in 19 MRs, they were neutral; and in the remaining 27, they disagreed or strongly disagreed.
Table 6: Responses to the MRs survey, for each statement (S1, S2, S3) in Table 3.
strongly agree agree neutral disagree strongly disagree
S1 3 46 12 3 0
S2 0 41 10 12 1
S3 0 18 19 24 3
@@
Regarding the non-positive responses to statement S3, we identified possible reasons from the practitioners’ qualitative feedback. In particular, one practitioner, who provided 11 ratings as “disagree” and one as “strongly disagree” to the statement, stated that while most of the MRs are correct and useful for reminding someone what to test, they are too generic to aid in deriving appropriate testing. This aligns with feedback from another practitioner at SISW who, despite not assessing the MRs, collected and reviewed the responses to the MRs questionnaire. The practitioner provided a possible reason for some of the MRs being very generic, stating that it could be attributed to the nature of the documents provided by SISW.
@@
Regarding the survey results for the 50 EMRs in EXP1, one practitioner provided consistent responses across all EMRs, stating they perceived that the EMRs were clean and understandable and consistent with their corresponding MRs. However, the practitioner deemed that not all functions used in the EMRs were feasible to be implemented. Regarding the last opinion, the practitioner noted that the functions in the EMRs, generated by ChatGPT, were specific to each EMR. Therefore, there is little opportunity for reusing these functions across different EMRs, meaning the effort required for implementation would be significant. In addition, the specific constraints of the testing environment for their applications may pose challenges in using the exposed APIs to implement functions in the EMRs. Even though EXP1 generated incomplete EMRs with regard to using the SUT’s APIs, as discussed above, this finding suggests that the characteristics of the SUT, such as its APIs, should be considered when further elaborating our approach to deriving EMRs.
The results for EXP1 indicate that our approach generates MRs and EMRs that practitioners deem to be understandable and relevant for testing.
@@
EXP2. In EXP2, we annotated each line of the ten EMRs, which were converted from the corresponding ten MRs studied in prior work [5]. These EMRs have a minimum of 10 statements, a mean of 13.6 statements, and a maximum of 20 statements, totaling 136 statements.
Table 7: Distribution of annotation labels (from Table 5) for the 10 EMRs obtained from EXP2.
C CLC AI WS WI IE INE ITE ES ENO WAU WLC MISS
52 54 1 3 0 3 1 9 1 2 0 10 3
38.2% 39.7% 0.7% 2.2% 0.0% 0.7% 0.0% 6.6% 0.7% 1.5% 0.0% 7.4% 2.2%
@@ C CLC AI WS WI IE INE ITE ES ENO WAU WLC MISS
Table 7 shows the results of EXP2 annotations using the labels defined in Table 5. Out of the 136 statements in the EMR code, the majority (107 statements, 78.6%) were classified as correctly converted. Specifically, 52 of the 107 statements are complex, and correctly and exactly implement part of the corresponding MRs, as shown by column C in Table 7. 54 of the 107 statements are simple, and correctly use SMRL constructs (column CLC). The remaining complex statement is valid and correct with respect to the corresponding MR (column AI). As for the 30 statements that were deemed incorrect, they can be further categorized as follows:
  • 16 statements incorrectly use APIs (labeled with IE, INE, ITE, ES, and ENO). For example, five ITE cases introduce the method isAuthorized to check if a user is authorized to perform an action, instead of verifying that the outputs obtained by the source and the follow-up inputs differ.
  • 10 statements incorrectly use SMRL constructs (column WLC). They are all related to the LLM “forgetting” how to use the construct IMPLIES; indeed, instead of separating the right-hand side of the implication with a comma (see end of Line 9 in Fig. 4), it uses an &. They can be easily fixed manually.
  • Three statements are wrong (column WS); for example, the generated MR checks if the returned page is an error page instead of the opposite.
  • Three statements miss implementing what is described in the corresponding explanations (column MISS); specifically, they do not compare the outputs of the source and the follow-up inputs. Further, these three statements had been annotated also with another label because, in addition to miss some instructions, they contained an incorrect implementation (i.e., ENO and WS), which is the reason why our labels sum up to 139 instead of 136.
The results for EXP2 indicate that the generated EMRs are largely correct with respect to their MRs.
5 Threats to Validity and Research Opportunities
Even though the experiment results are promising, we recognize that our work is in its preliminary stages, and there are some potential threats to its validity. In this section, we discuss these concerns, focusing on those that lead to challenging yet important directions for future research.
Prompt engineering. Our prompts are designed by decomposing the processes of deriving MRs from requirements and transforming MRs into EMRs into smaller, simpler steps. Despite these decomposed steps aligning with those that engineers would conduct manually, when using LLMs there might be better sequences of prompts that could improve the accuracy and relevance of the automatically generated MRs and EMRs. Further research is necessary to refine and compare the prompts, ensuring that they efficiently guide LLMs in producing effective MRs and EMRs.
Assessing MRs and EMRs. We evaluated the MRs and EMRs generated by our approach using the questionnaire-base survey and annotation studies, which rely on the quality of human assessments. Hence, such assessments are naturally biased by the participating personnel in the study. To minimize any threat to construct validity (i.e., researcher expectations affecting human subjects’ assessment), our industrial partner, SISW, invited four practitioners with different backgrounds for the questionnaire-based survey study. For the annotation study, the second author, who is familiar with SMRL, annotated the EMRs to ensure that the annotations are accurate. Although our choice may introduce a construct validity threat, the availability of experimental data enables other researchers to further investigate our findings.
As part of our future work, to better address face validity (i.e., the selection of appropriate reflective indicators), we aim to leverage recent studies on the assessment of code generation models [8] and select additional quantitative metrics to measure the degree to which MRs and EMRs are accurate and relevant to the SUT. Such metrics are desirable not only for objective assessment but also for automating the MT process; indeed, such metrics might be used as reward to improve LLMs results.
Human expertise. In our work, we leveraged human expertise to select technical documents and assess the outputs, i.e., MRs and EMRs. Since our approach does not involve humans-in-the-loop (e.g., the textual MRs generated by the LLM for EXP1 are not rewritten by humans before generating EMRs), the quality of outputs highly depends on the quality of inputs. To mitigate any bias originated from the input documents (construct validity), SISW selected four technical documents independently from the authors who developed the approach.
In the future, even if we aim to automate the MT process, incorporating humans in the loop could be use used to refine and validate the inputs and the intermediate outputs, to efficiently and effectively guide our approach to produce more accurate and relevant MRs and EMRs. Therefore, future research should explore methods to effectively incorporate human expertise into an AI-based MT process, balancing automation with human interventions.
External validity. To mitigate generalizability threats, for EXP1, we considered specifications of industrial modelling and simulation software for four different applications domains (marine design, wind turbine, air-craft propulsion, and aero-acoustics). Although such software may differ from other types of software systems (e.g., content management software), it is of high complexity and adopted in many industrial sectors including automotive and space. Further, for EXP2, we focused on MRs that demonstrated effective for content management (e.g., Joomla [11]) and Web-based automation software (e.g., Jenkins [7]), thus complementing the choice made for EXP1.
6 Related Work
Many MT approaches have been developed for testing various SUTs, such as search engines [26], Web applications [5], and embedded systems [1], accounting for the requirements specific to these SUTs. However, these approaches heavily rely on manual efforts to elicit MRs, inherently limiting the applicability of MT. To reduce the cost of defining MRs, researchers have proposed relying on meta-heuristic search [24, 23] and genetic programming [2] to automatically derive MRs from execution data. Although promising, such approaches require several executions of the SUT, which makes them suitable for small programs, but may not scale for large software where executions may take several minutes and a lot of outputs are produced.
Recently, a few studies [12, 25] have proposed the use of LLMs to automatically derive MRs from the knowledge base of LLMs. However, these studies are limited to deriving MRs for SUTs already known in the adopted LLMs. In other words, there are no studies demonstrating the applicability of LLMs to derive MRs that account for the requirements specific to an SUT that was unseen during the LLM training phases. Considering that, in industrial contexts, software testing activities often target new products implementing requirements different from those implemented by existing systems, LLM-based approaches for the automated generation of MRs should be capable of handling unseen requirements. This capability is what we presented in this paper.
Additionally, we note that, to the best of our knowledge, there is no work that automates the conversion of MRs into an executable form, which is required to fully automate the MT process. Chaleshtari et al. [5] proposed the adoption of a DSL to automate the execution of MRs; however, DSL-based MRs may have limited readability, contrary to MRs in natural language.
In summary, there is no solution in the literature that aims at fully automating the MT process, involving the derivation of MRs from SUT-specific requirements and the generation of EMRs, for the automated determination of test results in MT. An alternative is the automated generation of test cases—including oracles—using LLMs, which is under active development [22]. However, since LLM-generated code may still require human validation, we believe that validating EMRs is more cost-effective than validating test cases. Indeed, a single validated EMR enables exercising the SUT with multiple inputs, which might otherwise be exercised by different test cases, each needing validation.
7 Conclusion and Future work
@@
In this paper, we have introduced our approach for automatically deriving EMRs from requirements using LLMs. In our preliminary experiments, we applied our approach to four applications of our industry partner (SISW) and conducted a questionnaire-based survey study to collect opinions from SISW practitioners. In addition, we evaluated the correctness of the EMRs generated by our approach, by applying it to a Web application. The feedback from practitioners at SISW confirmed that our work is a promising direction for automating the derivation of MRs and EMRs from requirements. Furthermore, our annotation results indicate that our approach has strong potential for correctly generating EMRs from MRs. However, our results also reveal that the derived MRs need to be more specific to the SUTs, and the derived EMRs could be further improved with regard to the use of DSL constructs and APIs, which necessitate further research.
@@
In our future work, we plan to explore several key areas related to our current research. A primary focus will be on automating prompt engineering to facilitate the generation of prompts for deriving MRs from requirements and their subsequent conversion into EMRs. Recognizing the significance of input quality, particularly in requirements, another important area will involve developing methods to assist engineers in providing high-quality inputs. Integrating human expertise, or “human in the loop”, will also be an important aspect of our research, aiming to further enhance the accuracy and relevance of the generated MRs and EMRs.